OpenStudy (anonymous):

How would I solve this equation for the variable y? m = 2/5 y + n

geerky42 (geerky42):

try to isolate y. can you do it?

OpenStudy (anonymous):

I think I would subtract n first? I'm not sure though..

geerky42 (geerky42):

exactly. do it, then you got m - n = 2/5 y y is still not isolated yet. do you know what to do?

OpenStudy (anonymous):

I know I need to somehow get rid of the 2/5, but I don't know how I would do that.

geerky42 (geerky42):

first, we get rid of denominator. just multiply both sides by 5. \[m - n = \frac{2}{5}y \rightarrow 5(m- n)=2y \] is this clear? do you know what to do next?

OpenStudy (anonymous):

Okay. So it would then be 5 (m - n) = 2y And I would divide both sides by two..?

geerky42 (geerky42):

correct then simply left side of the equation does this make sense?

OpenStudy (anonymous):

How could the left side be simplified?

OpenStudy (anonymous):

I understand it up to that point.

geerky42 (geerky42):

i mean simply this: \[\frac{5(m-n)}{2}\]

OpenStudy (anonymous):

Ohh okay. Misread it. xD Thank you so much for your help!

geerky42 (geerky42):

glad i helped :)